the speed of a stream is 3 km/hr. a boat travels 4 km upstream in the same time it takes to travel 10 km downstream. What is the speed of the boat in still water?
Answer by aswathytony(47)

let the speed of boat in still water is x km/hr.
given speed of stream is 3 km/hr.
then speed of boat upstream = ( x - 3) km/hr
speed of boat downstream = (x + 3 ) km/hr
Distance travelled upstream = 4 km
time taken to travel upstream = distance travelled upstream / speed of boat
upstream
= 4 / (x - 3 ) .............(1)

Distance travelled downstream = 10 km
time taken to travel downstream = distance travelled downstream / speed of boat
downstream
= 10 / (x + 3) ............(2)
given that boat takes same time to travel upstream and downstream.i.e; (1) & (2) are equal
i.e. 4/(x-3) = 10/ (x + 3 )
4 * (x +3 ) = 10 * (x- 3)
4x + 12 = 10x - 30
12 + 30 = 10x -4x
42 = 6x
x = 42 / 6 = 7.
i.e speed of boat in still water is 7 km/hr.